OpenStudy (anonymous):

given thatp(a)=0.4, and p(b)=0.5, and p(a andB)=0.16, determine p(bla)

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\]

OpenStudy (anonymous):

oops you want \[P(B|A)=\frac{P(A\cap B)}{P(A)}\]

OpenStudy (anonymous):

so \(P(B|A)=\frac{.16}{.4}=\frac{1.6}{4}=.4\)